# Doomed Worlds I: No new evidence for orbital decay in a long-term survey of 43 ultra-hot Jupiters

### v1

@@ 2.4 Minerva-Australis observations and photometry
An instrumental glitch in the middle of the sequence caused a jump with image rotation, followed by slow de-rotation back to the original alignment, which combined with deteriorating weather resulted in poorer data quality after midtransit. To control for the effect of the rotation, the final light curve was created using a separate Mathematica photometry notebook to do differential circular-aperture photometry on the target star and two nearby, bright comparisons. Stars were centroided on every frame to determine the centers of the apertures. For each star, three hand-selected boxes well outside of the apertures and without stars were used to determine the background value. The final light curve was generated by subtracting the background counts from each star, dividing the target by the average signal of the comparison stars, and normalizing to the median out-of-transit baseline. The lowest baseline scatter was achieved with an aperture 21 pixels in diameter.
@@ 2.5 IRTF and Las Campanas Observatory observations and photometry
Three transits of CoRoT-2 b were observed in 2008 shortly after the planet’s discovery was announced, and they were included in Adams (2010) but the light curves have not been published until now. Two transits were taken on the 6.5 m Magellan Clay telescope at Las Campanas Observatory (LCO) on 2008 Apr 24 and 2008 May 15 during the first observing runs to use the updated MagIC-e2v camera (see e.g., Osip et al., 2004; Adams et al., 2011). MagIC-e2v had 1024x1024 pixels, each 13 micron square, and on Magellan had a field of view of 38″ x 38″ and a plate scale of 0.037″ per pixel unbinned. The timing for these observations came from network computers that were verified each evening to be synced within a second to UTC standard. Note that the partial transit on 2008 Apr 24 was deliberately scheduled right before dawn after two other transits were observed that night and is not due to weather.
The third transit of CoRoT-2 b on 2008 Sep 7 was observed using the MIT Optical Rapid Imaging System (MORIS) on NASA’s 3-m Infrared Telescope Facility (IRTF; Gulbis et al., 2011). A dichroic was used to direct light to MORIS, through a Thor long-pass red visitor filter with a lower cutoff at 700 nm. The plate scale for MORIS is 0.114″ per pixel, with a 60″ x 60″ field of view. We used the 1 MHz Conventional mode in 2.4x gain setting, which has a read noise of 6 e- per pixel and gain of 1.5e-/ADU. MORIS observations were triggered by GPS and had better than microsec timing accuracy.
Additionally, we observed two transits of TrES-1b using MORIS on the IRTF on 2009 June 28 and 2009 Jul 1, using the same settings as for CoRoT-2 b, which have not been previously published. Observational details for all of the IRTF and LCO transits are included in Table 1.
@@ 3 Data from Space Telescopes
In addition to new ground-based observations, we present new fit results to light curves from at least one TESS sector for all six planets, and new fits for the CoRoT mission data for CoRoT-2 b. Although all space-based data used to make these light curves is publicly available and should remain that way in perpetuity, the version of the data available to the public may change in the future (as happened for the CoRoT data). The TESS data also required non-trivial processing (smoothing and detrending) to produce these light curves. For these reasons, we consider it important to publish the version of the transit light curves that we fit in this work, so as to make it as easy as possible to reproduce our results and/or to identify the source of problems that may come to light in the future. All transit light curves may be found in a supplementary electronic file, with a stub table for format at Table 2.
@@ 3.2 CoRoT photometry
For CoRoT-2 b, we downloaded data from the LEGACY data release (version 4)222 http://idoc-corot.ias.u-psud.fr/sitools/client-user/COROT_N2_PUBLIC_DATA/project-index.html. We used the red flux (keyword “REDFLUX”), which was less noisy than either the green or the blue fluxes, and the frame times given in BJD/TT (keyword “DATEBARTT”); for an explanation of why the ms-level difference between TT (Terrestrial Time) and TDB (Barycentric Dynamical Time) is not important for our purposes, see the discussion of timing systems in Section 5.3. For CoRoT-2 b, the first three transits (out of 82) in the 150-day sequence were recorded at the longer cadence (512 s), at which point the transit signature was identified by the CoRoT team, and the rest of the data were taken at 32 s. We chose not to smooth and flatten the CoRoT data in the same way as we did the TESS data, but instead to use a subset of the data in an eight-hour window centered around the predicted transit midtime. The linear term in our fit function accounted for any remaining trends, which were slight.
| Value Used | Observations | |||||||||
| Planet | P (d) | a/ | i | Age (Gyr) | Epochs | Years | ||||
@@
| From the Exoplanet Characterization Catalog (Kokori et al., 2022) and NASA Exoplanet Archive. | ||||||||||
| Number of transit midtimes used in this work. | ||||||||||
| Number of epochs between first and last transit midtime. | ||||||||||
| Years spanned by observed transit midtimes. | ||||||||||
@@ 4 Transit light-curve fitting
Each transit light curve was fit using Markov chain Monte Carlo (MCMC) methods using the publicly available python package PyLightcurve (Tsiaras et al., 2016), which is based on Emcee, an affine-invariant ensemble sampler (Foreman-Mackey et al., 2013) and makes use of several other scientific packages and catalogs333 matplotlib: (Hunter, 2007); ExoTETHyS: (Morello et al., 2020); Exoplanet Characterisation Catalogue (Kokori et al., 2022); astropy (Astropy Collaboration et al., 2013); scipy: (Virtanen et al., 2020); numpy: https://archive.org/details/NumPyBook. For each fit we used 400 walkers and 150,000 links, with the first 30,000 links as burn-in. Convergence of chains was checked by examining the autocorrelation time as well as examining the traces and the correlations between parameters for signs of non-convergence.
Our focus for this work is on the timing, and most of the transit light curve parameters were fixed to the best-fit parameters from more precise light curves, including the eccentricity (), inclination (, and the ratio of the orbital semimajor axis to the stellar radius (). We used the values reported in Table 3 taken from the Exoplanet Characterization Catalog, which was developed for the ExoClock project (Kokori et al., 2022), with missing or (rarely) inaccurate parameters replaced by values from the the NASA Exoplanet Archive where necessary. Note that for most light curves in this work the signal to noise ratio is modest enough that the fitted transit midtime values are quite insensitive to the choice of values for, e.g., or the limb darkening coefficients. We nonetheless verified that the fits were reasonable by checking for abnormal scatter in the residuals from the model fits, especially around ingress and egress where issues with the shape. Limb-darkening coefficients were calculated for the appropriate bandpass in pylightcuve using ExoTethys (Morello et al., 2020). For the six planets that are the focus of this work – CoRoT-2 b, TrES-1 b, WASP-12 b, WASP-19 b, WASP-46 b, and WASP-121 b – we have published all parameters, fitted and fixed, for each light curve that we fit, including those from CoRoT and TESS, which is available as a supplementary table online with a stub table for format at Table 4. We are also publishing the transit midtimes and errors as a standalone table, along with their timing systems and sources, to aid future timing analysis efforts (Table 5).
@@
| Parameter | Value | Err minus | Err plus |
@@ 5 Curation of Literature Light-Curve Times
Although regular, new observations are key to a long-term project such as this one, it is the entire cumulative body of transit knowledge that allows us to make sensitive timing measurements. For several targets, the literature data now span two decades, and some planets have 10+ reference works containing hundreds of transits observed on many telescopes using a variety of methods to generate and fit the transit light curves. Since most of these works contain a mix of new data, reanalyzed data, and times drawn directly or with some correction from the literature, data curation has become and will remain a large component of precise-timing analyses. We discuss below some of the lessons learned, including the importance of identifying duplicated times (Section 5.1) and composite times (Section 5.2), and the ever-present pitfall of improper conversion from the JD/UTC and BJD/TDB timing systems (Section 5.3).
@@
As our starting point for compiling the literature midtimes, we used the work of Ivshina & Winn (2022), which contains both new analyses of TESS light curves and a thorough literature review through 2022 for hundreds of hot Jupiters. Although we have identified some errors in Ivshina & Winn (2022) that are discussed below, as is perhaps inevitable for any static database of this magnitude, we note that work represents a formidable contribution to the timing literature. In particular, the table in Ivshina & Winn (2022) has two extremely useful features for error corrections: (1) it converts the literature times into the BJD/TDB time system wherever possible, and explicitly states the time system used, and (2) it provides the Astrophysics Data System (ADS) code for the paper where each transit midtime was reported, making it easier to identify the source of every reported midtime. Without these features it would have been much harder to uncover the errors that we have found.
@@
Our second main source of transit midtimes is the Exoplanet Transit Database, or ETD444http://var2.astro.cz/ETD/, which provides an online repository for light curves from amateur astronomers, who have posted tens of thousands of light curves of hundreds of systems since the website went online in September 2008 (Poddaný et al., 2010). Although the data on the ETD can vary widely in quality, many planets have been observed with nearly continuous timing coverage, a key feature for constraining long-term timing variations. For each planet, the ETD provides a table listing all of their available midtimes, which we used with three additional cuts to the data: (1) for systems with many transits, we selected their highest quality data flag (DQ=1 for CoRoT-2 b, TrES-1b, WASP-12 b, and WASP-19 b, and DQ=1 or 2 for WASP-46 b and WASP-121 b); (2) we removed all transits with very large midtime errors ( min); and (3) we removed duplicate times that are in the ETD but are also in Ivshina & Winn (2022), preferring the values from Ivshina & Winn (2022), which are in the BJD/TDB system (the ETD uses HJD/UT; see discussion of timing systems in Section 5.3). Looking for duplicate transits had to be done by hand to ensure that we did not inadvertently remove points that were observed on the same night but at two different sites, and will now be discussed in more detail.
@@ 5.1 Multiple reported transit midtimes for the same epoch
When results are combined from many sources, it is common for there to be multiple, non-identical transit midtimes reported at the same transit epoch. The challenge is to identify which are duplicates that need to be weeded out, and which are simultaneous observations that should be kept. We have found at least four ways in which repeats occur: (1) two or more telescopes were used to observe during the same transit epoch; (2) a single telescope simultaneously observed a transit in two or more wavelengths, producing multiple transit light curves; (3) two or more groups used the same observations to generate different transit light curves and/or made different model fits; and (4) a transit midtime was reported in one timing system (e.g., HJD/UTC) in one work and then was republished later in another timing system (e.g., BJD/TDB). In the first two cases, all of the reported midtimes are independent observations and should be included in timing analyses, but in cases 3 and 4 only one midtime value should be used. Special care needs to be taken with systems that have been repeatedly analyzed, since compilations of compilations often obscure both the source and the original value of the reported transit midtimes.
@@
For WASP-43 b, each transit from Wang et al. (2021) that also appeared in Patra et al. (2020) and either Gillon et al. (2012) or Murgas et al. (2014) shows up twice in Ivshina & Winn (2022). Since the duplicated transits pre-date any of the new observations listed in Table 1 in Wang et al. (2021), the listed times must have been sourced from one of the two prior works. Removing the 24 duplicate times from Wang et al. (2021) caused the value to fall in half, and the best fit curvature went from a tiny but 3 positive trend with ms yr to less than 1 from zero (see Table 7).
@@
Given the large number of works involved, it is likely there are other instances of duplication in Ivshina & Winn (2022), especially when citing works such as Wang et al. (2021) which do not list the sources of individual transit midtimes. It is important to note that while impact on of of including duplicates is often small, it can be substantial if the duplicated point has very low errors or occurs at a critical point in the sequence. Duplications are are also an entirely unnecessary source of error that could be eliminated with better data management practices by the field as a whole. It is recommended that researchers carefully investigate the source of each transit time used in future analyses, particularly before making claims of subtle detections of small effects. For a long-term approach to how the transit timing community could deal with this problem, see our recommendations in Section 8.3.
@@ 5.2 Composite transit times
We now turn to the related issue of composite transit times, or times that are derived from more than one transit light curve. We have identified at least three potential issues that may arise when attempting to use times derived from more than one transit. (1) Composite times typically have error bars that are much smaller than any individual transit, and they compress all of the timing information into a single point, which can distort analyses of long-term variations. (2) Composite times obscure the source of the original timing and make error detection and correction difficult. And (3) composite times may lead to double-counting transits, if one or more of the individual transits that were used to calculate the composite time were also published on their own.
@@
Stacked transits result from combining photometric time-series data spanning multiple transit epochs into a single transit light curve. Long photometric time series often have a low signal-to-noise ratio or low sampling for individual transits, and it may not be possible to fit, or even detect, individual transits. By folding on the best-fit orbital period and stacking the data, it is possible to extract a single, higher-quality light curve. Stacked transits are common in discovery papers from large ground-based surveys, as well as for detecting very small planets in Kepler/K2 data (e.g., Adams et al., 2021). Kepler-1658 b, for instance, is only robustly detected by stacking all transits within an individual Kepler quarter or TESS sector (Vissapragada et al., 2022). In principle, there is no issue with using a midtime that was fit to a stacked transit, provided the data being stacked came from a single origin (e.g, the same ground-based survey) and have roughly consistent noise levels and systematics. It is common practice to assign the midtime of the stacked transit to either the first transit epoch in the sequence or to the transit epoch nearest to the sequence midpoint. Often, though, particularly for survey discovery data, the assigned midtime for a stacked transits may not be published, except as it may contribute to the fitted value for an ephemeris midtime (), discussed next.
@@
Ephemeris midtimes, or , are the result of fitting a set of individual transit midtimes to the linear ephemeris in Equation 2 and may be called variously the reference midtime, the time of conjunction, the time of transit, or similar. In some cases, is the only timing information provided in a published work. Although it should be common practice to publish the photometric time series for each transit light curve, many papers still do not do so. Nor do papers always report fitted midtimes for each transit light curve. For systems with sparse data, an ephemeris midtime may be one of the few timing points available, and frequently the midtime at the very first epoch is actually an ephemeris midtime. However, using a value in combination with midtimes derived from individual transits can be problematic. A value based on multiple transits taken over many months or years will have a much smaller error bar than any of its component transits, or indeed any individual transit, and thus will serve to anchor the model to the assigned transit epoch. In many cases, this results in a larger value for than would result if the individual transit midtimes and errors had been used instead.
@@
The third type of composite time is a weighted-average time, where two or more midtimes from independent observations (say, transits at two different telescopes) are averaged together. We do not recommend publishing or using weighted-average times, since timing models can easily deal with multiple data points at the same epoch. Moreover, it is impossible to reconstruct an individual midtime and error from a weighted average, which also makes it impossible to assess for mistakes in timing system conversion, or sometimes even to determine which transits were used to calculate the average time.

### v2

@@ 2.4 Minerva-Australis observations and photometry
An instrumental glitch in the middle of the sequence caused a jump with image rotation, followed by slow derotation back to the original alignment, which combined with deteriorating weather resulted in poorer data quality after midtransit. To control for the effect of the rotation, the final light curve was created using a separate Mathematica photometry notebook to do differential circular-aperture photometry on the target star and two nearby, bright comparisons. Stars were centroided on every frame to determine the centers of the apertures. For each star, three hand-selected boxes well outside of the apertures and without stars were used to determine the background value. The final light curve was generated by subtracting the background counts from each star, dividing the target by the average signal of the comparison stars, and normalizing to the median out-of-transit baseline. The lowest baseline scatter was achieved with an aperture 21 pixels in diameter.
@@ 2.5 IRTF and Las Campanas Observatory Observations and Photometry
Three transits of CoRoT-2 b were observed in 2008 shortly after the planet’s discovery was announced, and they were included in Adams (2010) but the light curves have not been published until now. Two transits were taken on the 6.5 m Magellan Clay telescope at Las Campanas Observatory (LCO) on 2008 April 24 and 2008 May 15 during the first observing runs to use the updated MagIC-e2v camera (see e.g., Osip et al., 2004; Adams et al., 2011). MagIC-e2v had 1024x1024 pixels, each 13 micron square, and on Magellan had a field of view of 38″ x 38″ and a plate scale of 0.037″ per pixel unbinned. The timing for these observations came from network computers that were verified each evening to be synced within a second to UTC standard. Note that the partial transit on 2008 Apr 24 was deliberately scheduled right before dawn after two other transits were observed that night and is not due to weather.
The third transit of CoRoT-2 b on 2008 September 7 was observed using the MIT Optical Rapid Imaging System (MORIS) on NASA’s 3-m Infrared Telescope Facility (IRTF; Gulbis et al., 2011). A dichroic was used to direct light to MORIS through a Thor long-pass red visitor filter with a lower cutoff at 700 nm. The plate scale for MORIS is 0.114″ per pixel, with a 60″ x 60″ field of view. We used the 1 MHz Conventional mode in 2.4x gain setting, which has a read noise of 6 e- per pixel and gain of 1.5e-/ADU. MORIS observations were triggered by GPS and had better than microsecond timing accuracy.
Additionally, we observed two transits of TrES-1 b using MORIS on the IRTF on 2009 June 28 and 2009 July 1, using the same settings as for CoRoT-2 b, which have not been previously published. Observational details for all of the IRTF and LCO transits are included in Table 1.
@@ 3 Data from Space Telescopes
In addition to new ground-based observations, we present new fit results to light curves from at least one TESS sector for all six planets and new fits for the CoRoT mission data for CoRoT-2 b. Although all space-based data used to make these light curves is publicly available and should remain that way in perpetuity, the version of the data available to the public may change in the future (as happened for the CoRoT data). The TESS data also required nontrivial processing (smoothing and detrending) to produce these light curves. For these reasons, we consider it important to publish the version of the transit light curves that we fit in this work, so as to make it as easy as possible to reproduce our results and/or to identify the source of problems that may come to light in the future. All transit light curves may be found in a supplementary electronic file, with a stub table for format at Table 2.
@@ 3.2 CoRoT photometry
For CoRoT-2 b, we downloaded data from the LEGACY data release (version 4)222 http://idoc-corot.ias.u-psud.fr/sitools/client-user/COROT_N2_PUBLIC_DATA/project-index.html. We used the red flux (keyword “REDFLUX”), which was less noisy than either the green or the blue fluxes, and the frame times given in BJD/TT (keyword “DATEBARTT”); for an explanation of why the ms-level difference between TT (Terrestrial Time) and TDB (Barycentric Dynamical Time) is not important for our purposes, see the discussion of timing systems in Section 5.3. For CoRoT-2 b, the first three transits (out of 82) in the 150-day sequence were recorded at the longer cadence (512 s), at which point the transit signature was identified by the CoRoT team, and the rest of the data were taken at 32 s. We chose not to smooth and flatten the CoRoT data in the same way as we did the TESS data but instead use a subset of the data in an eight-hour window centered around the predicted transit midtime. The linear term in our fit function accounted for any remaining trends, which were slight.
| Value Useda | Observations | |||||||||
| Planet | P (d) | a/ | i | Age (Gyr) | b | Epochsc | Yearsd | |||
@@
| aFrom the Exoplanet Characterization Catalog (Kokori et al., 2022) and NASA Exoplanet Archive. | ||||||||||
| bNumber of transit midtimes used in this work. | ||||||||||
| cNumber of epochs between first and last transit midtime. | ||||||||||
| dYears spanned by observed transit midtimes. | ||||||||||
@@ 4 Transit light-curve fitting
Each transit light curve was fit using Markov Chain Monte Carlo (MCMC) methods using the publicly available python package PyLightcurve (Tsiaras et al., 2016), which is based on Emcee, an affine-invariant ensemble sampler (Foreman-Mackey et al., 2013) and makes use of several other scientific packages and catalogs333 matplotlib: (Hunter, 2007); ExoTETHyS: (Morello et al., 2020); Exoplanet Characterisation Catalogue (Kokori et al., 2022); astropy (Astropy Collaboration et al., 2013); scipy: (Virtanen et al., 2020); numpy: https://archive.org/details/NumPyBook. For each fit we used 400 walkers and 150,000 links, with the first 30,000 links as burn-in. Convergence of chains was checked by examining the autocorrelation time as well as the traces and the correlations between parameters for signs of nonconvergence.
Our focus for this work is on the timing, and most of the transit light curve parameters were fixed to the best-fit parameters from more precise light curves, including the eccentricity (), the inclination (, and the ratio of the orbital semimajor axis to the stellar radius (). We used the values reported in Table 3 taken from the Exoplanet Characterization Catalog, which was developed for the ExoClock project (Kokori et al., 2022), with missing or (rarely) inaccurate parameters replaced by values from the the NASA Exoplanet Archive where necessary. Note that for most light curves in this work the signal-to-noise ratio is modest enough that the fitted transit midtime values are quite insensitive to the choice of values for, e.g., or the limb darkening coefficients. We nonetheless verified that the fits were reasonable by checking for abnormal scatter in the residuals from the model fits, especially around ingress and egress where issues with the shape are most evident. Limb-darkening coefficients were calculated for the appropriate bandpass in pylightcuve using ExoTethys (Morello et al., 2020). For the six planets that are the focus of this work – CoRoT-2 b, TrES-1 b, WASP-12 b, WASP-19 b, WASP-46 b, and WASP-121 b – we have published all parameters, fitted and fixed, for each light curve that we fit, including those from CoRoT and TESS, which are available as a supplementary table with a stub table for format at Table 4. We are also publishing the transit midtimes and errors as a standalone table, along with their timing systems and sources, to aid future timing analysis efforts (Table 5).
@@
| Parameter | Value | Err minusc | Err plus |
@@ 5 Curation of Literature Light-Curve Times
Although regular new observations are key to a long-term project such as this one, it is the entire cumulative body of transit knowledge that allows us to make sensitive timing measurements. For several targets, the literature data now span two decades, and some planets have 10+ reference works containing hundreds of transits observed on many telescopes using a variety of methods to generate and fit the transit light curves. Since most of these works contain a mix of new data, reanalyzed data, and times drawn directly or with some correction from the literature, data curation has become and will remain a large component of precise timing analyses. We discuss below some of the lessons learned, including the importance of identifying duplicated times (Section 5.1) and composite times (Section 5.2), and the ever-present pitfall of improper conversion from the JD/UTC and BJD/TDB timing systems (Section 5.3).
@@
As our starting point for compiling the literature midtimes, we used the work of Ivshina & Winn (2022), which contains both new analyses of TESS light curves and a thorough literature review through 2022 for hundreds of hot Jupiters. Although we have identified some errors in Ivshina & Winn (2022) that are discussed below, as is perhaps inevitable for any static database of this magnitude, we note that their work represents a formidable contribution to the timing literature. In particular, the table in Ivshina & Winn (2022) has two extremely useful features for error corrections: (1) it converts the literature times into the BJD/TDB time system wherever possible and explicitly states the time system used, and (2) it provides the Astrophysics Data System (ADS) code for the paper where each transit midtime was reported, making it easier to identify the source of every reported midtime. Without these features it would have been much harder to uncover the errors that we have found.
@@
Our second main source of transit midtimes is the Exoplanet Transit Database, or ETD444http://var2.astro.cz/ETD/, which provides an online repository for light curves from amateur astronomers, who have posted tens of thousands of light curves of hundreds of systems since the website went online in September 2008 (Poddaný et al., 2010). Although the data on the ETD can vary widely in quality, many planets have been observed with nearly continuous timing coverage, a key feature for constraining long-term timing variations. For each planet, the ETD provides a table listing all of their available midtimes, which we used with three additional cuts to the data: (1) for systems with many transits, we selected their highest quality data flag (DQ=1 for CoRoT-2 b, TrES-1b, WASP-12 b, and WASP-19 b, and DQ=1 or 2 for WASP-46 b and WASP-121 b); (2) we removed all transits with very large midtime errors ( minutes); and (3) we removed duplicate times that are in the ETD but are also in Ivshina & Winn (2022), preferring the values from Ivshina & Winn (2022), which are in the BJD/TDB system (the ETD uses HJD/UT; see discussion of timing systems in Section 5.3). Looking for duplicate transits had to be done by hand to ensure that we did not inadvertently remove points that were observed on the same night but at two different sites and will now be discussed in more detail.
@@ 5.1 Multiple reported transit midtimes for the same epoch
When results are combined from many sources, it is common for there to be multiple, nonidentical transit midtimes reported at the same transit epoch. The challenge is to identify which are duplicates that need to be weeded out and which are simultaneous observations that should be kept. We have found at least four ways in which repeats occur: (1) two or more telescopes were used to observe during the same transit epoch; (2) a single telescope simultaneously observed a transit in two or more wavelengths, producing multiple transit light curves; (3) two or more groups used the same observations to generate different transit light curves and/or made different model fits; and (4) a transit midtime was reported in one timing system (e.g., HJD/UTC) in one work and then was republished later in another timing system (e.g., BJD/TDB). In the first two cases, all of the reported midtimes are independent observations and should be included in timing analyses, but in cases 3 and 4 only one midtime value should be used. Special care needs to be taken with systems that have been repeatedly analyzed, since compilations of compilations often obscure both the source and the original value of the reported transit midtimes.
@@
For WASP-43 b, each transit from Wang et al. (2021) that also appeared in Patra et al. (2020) and either Gillon et al. (2012) or Murgas et al. (2014) shows up twice in Ivshina & Winn (2022). Since the duplicated transits predate any of the new observations listed in Table 1 in Wang et al. (2021), the listed times must have been sourced from one of the two prior works. Removing the 24 duplicate times from Wang et al. (2021) caused the value to fall by half, and the best-fit curvature went from a tiny but 3 positive trend with ms yr-1 to less than 1 from zero (see Table 7).
@@
Given the large number of works involved, it is likely there are other instances of duplication in Ivshina & Winn (2022), especially when citing works such as Wang et al. (2021) which do not list the sources of individual transit midtimes. It is important to note that while the impact on of including duplicates is often small, it can be substantial if the duplicated point has very low errors or occurs at a critical point in the sequence. Duplications are also an entirely unnecessary source of error that could be eliminated with better data management practices by the field as a whole. It is recommended that researchers carefully investigate the source of each transit time used in future analyses, particularly before making claims of subtle detections of small effects. For a long-term approach to how the transit timing community could deal with this problem, see our recommendations in Section 8.3.
@@ 5.2 Composite transit times
We now turn to the related issue of composite transit times, or times that are derived from more than one transit light curve. We have identified at least three potential issues that may arise when attempting to use times derived from more than one transit. (1) Composite times typically have error bars that are much smaller than any individual transit, and they compress all of the timing information into a single point, which can distort analyses of long-term variations. (2) Composite times obscure the source of the original timing and make error detection and correction difficult. (3) Composite times may lead to double-counting transits, if one or more of the individual transits that were used to calculate the composite time were also published on their own.
@@
Stacked transits result from combining photometric time-series data spanning multiple transit epochs into a single transit light curve. Long photometric time series often have a low signal-to-noise ratio or low sampling for individual transits, and it may not be possible to fit, or even detect, individual transits. By folding on the best-fit orbital period and stacking the data, it is possible to extract a single, higher-quality light curve. Stacked transits are common in discovery papers from large ground-based surveys, as well as for detecting very small planets in Kepler/K2 data (e.g., Adams et al., 2021). Kepler-1658 b, for instance, is only robustly detected by stacking all transits within an individual Kepler quarter or TESS sector (Vissapragada et al., 2022). In principle, there is no issue with using a midtime that was fit to a stacked transit, provided the data being stacked came from a single origin (e.g, the same ground-based survey) and have roughly consistent noise levels and systematics. It is common practice to assign the midtime of the stacked transit to either the first transit epoch in the sequence or the transit epoch nearest to the sequence midpoint. Often, though, particularly for survey discovery data, the assigned midtime for a stacked transit may not be published, except as it may contribute to the fitted value for an ephemeris midtime (), discussed next.
@@
Ephemeris midtimes, or , are the result of fitting a set of individual transit midtimes to the linear ephemeris in Equation 2 and may be variously called the reference midtime, the time of conjunction, the time of transit, or similar. In some cases, is the only timing information provided in a published work. Although it should be common practice to publish the photometric time series for each transit light curve, many papers still do not do so. Nor do papers always report fitted midtimes for each transit light curve. For systems with sparse data, an ephemeris midtime may be one of the few timing points available, and frequently the midtime at the very first epoch is actually an ephemeris midtime. However, using a value in combination with midtimes derived from individual transits can be problematic. A value based on multiple transits taken over many months or years will have a much smaller error bar than any of its component transits, or indeed any individual transit, and thus will serve to anchor the model to the assigned transit epoch. In many cases, this results in a larger value for than would result if the individual transit midtimes and errors had been used instead.
@@
The third type of composite time is a weighted-average time, where two or more midtimes from independent observations (say, transits at two different telescopes) are averaged together. We do not recommend publishing or using weighted-average times, since timing models can easily deal with multiple data points at the same epoch. Moreover, it is impossible to reconstruct an individual midtime and error from a weighted average, which also makes it impossible to assess for mistakes in timing system conversion or sometimes even to determine which transits were used to calculate the average time.